# Matching with Generalized Sequential Reserves

### v2

@@ 4. Stable and Strategy-proof Mechanism Design
We now consider a centralized market in which there is a finite set of individuals $I=\{i_{1},...,i_{n}\}$ and a finite set of institutions $S=\{s_{1},...,s_{m}\}$. We let $\overline{q}_{s}$ be the capacity of the institution $s$. There is a finite set of contracts $X$. $\mathbf{i}(x)$ and $\mathbf{s}(x)$ denote the individual and institution associated with the contract $x\in X$, respectively.
@@
For any set of contracts $Y\subseteq X$, let $\mathbf{i}(Y)\equiv\bigcup_{x\in Y}\{\mathbf{i}(x)\}$ be the set of individuals who have a contract associated with them in $Y$. Similarly, we let $\mathbf{s}(Y)\equiv\bigcup_{x\in Y}\{\mathbf{s}(x)\}$. We also define $Y_{i}\equiv\{x\in Y\mid i=\mathbf{i}(x)\}$ for each $i\in I$ and $Y_{s}\equiv\{x\in Y\mid s=\mathbf{s}(x)\}$ for each $s\in S$.
@@
Each individual $i\in I$ has a strict preference $P_{i}$ over contracts in $X_{i}$ and an outside option $\emptyset$. A contract $x\in X_{i}$ is acceptable to $i$ with respect to $P_{i}$ if $xP_{i}\emptyset$. We let $R_{i}$ be the weak preference relation associated with $P_{i}$. $xR_{i}y$ if and only if $xP_{i}y$ or $x=y$.
@@
Each institution $s\in S$ has multiunit demand and is endowed with an aggregate choice rule $C_{s}$ that describes how $s$ would choose from any set of contracts. For all $Y\subseteq\ X$ and $s\in S$, the aggregate choice rule $C_{s}$ is such that $C_{s}(Y)\subseteq Y_{s}$ and $\mid C_{s}(Y)\mid\leq\overline{q}_{s}$.
A matching is a set of contracts $Y\subseteq X$. A matching $Y\subseteq X$ is feasible if $\mid Y_{i}\mid\leq 1$ for all $i\in I$. A feasible outcome $Y\subseteq X$ is stable if
@@
Condition (1) is the standard individual rationality condition for individuals. Condition (2) requires that institutions’ selection procedures are respected. This condition ensures that criteria, such as diversity and affirmative action, that are encoded in institutions’ choice rules are complied. [^4] See Alva and Dogan (2023) for a detailed discussion of this property. Condition (3) is the standard no blocking condition and requires that there be no set of contracts $Z$ such that all institutions and individuals associated with contracts in $Z$ prefer to receive the contracts in $Z$.
Stability and strategy-proofness are fundamental objectives in market design. Stability precludes scenarios where agents can profitably rearrange their matches after market-clearing. Strategy-proofness establishes truthful preference revelation as a dominant strategy; this eliminates incentives for strategic misrepresentation, thereby simplifying market participation and ensuring that allocations derive from accurate preference data.
Given a profile of aggregate choice rules of institutions $C=(C_{s})_{s\in\ S}$, a mechanism $\psi(\cdot;C)$ maps preference profiles $P=(P_{i})_{i\in\ I}$ to outcomes. Unless otherwise stated, we assume that institutions’ choice rules are fixed. A mechanism $\psi$ is stable if $\psi(P;C)$ is a stable outcome for every preference profile $P$, given a profile of the institutions’ choice rules $C$.
A mechanism $\psi$ is strategy-proof if, given a profile of institutions’ choice rules $C$, for every preference profile $P$ and for each individual $i\in I$, there is no $\widetilde{P}_{i}$ such that
@@
Hatfield and Kominers (2019) introduce conditions on the choice rules of the institutions to guarantee a stable and strategy-proof matching mechanism.
Definition 3 (Hatfield and Kominers (2019)).
@@
Let $C_{s}(\cdot;q_{s})$ be an GS choice rule induced by the transfer function $q_{s}$. We define a related choice function $\overline{C}_{s}(\cdot;q_{s})$ as follows:
@@
Given $Y^{1}$ and $\overline{q}^{1}_{s}$, division 1 chooses $\overline{C}^{1}_{s}(Y^{1};\overline{q}^{1}_{s})$. Let $r_{1}=\overline{q}^{1}_{s}-\mid\overline{C}^{1}_{s}(Y^{1};\overline{q}^{1}_{s})\mid$. Let $Y^{2}\equiv Y^{1}\setminus\overline{C}^{1}_{s}\left(Y^{1};\overline{q}^{1}_{s}\right)$.
@@
Given $Y^{k}$ and its capacity $q^{k}_{s}(r_{1},...,r_{k-1})$, division $k$ chooses $\overline{C}^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))$. Let $r_{k}=q^{k}_{s}(r_{1},...,r_{k-1})-\mid\overline{C}^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))\mid$. Let $Y^{k+1}\equiv Y^{k}\setminus\overline{C}^{k}_{s}\left(Y^{k};q^{k}_{s}\left(r_{1},\ldots,r_{k-1}\right)\right)$.
The union of divisions’ choices is the institution’s chosen set. That is,
$\overline{C}_{s}(Y;q_{s})\equiv\overline{C}^{1}_{s}(Y^{1};\overline{q}^{1}_{s})\cup\bigcup_{k=2}^{K_{s}}{\overline{C}^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))}.$
The difference between $C_{s}$ and $\overline{C}_{s}$ is as follows: In the computation of $C_{s}$, if a contract of an individual is chosen by some category, then her other contracts are removed for the rest of the procedure. According to $\overline{C}_{s}$, if an individual’s contract is chosen by division $k$, then her other contracts will still be available for the remaining divisions.
@@
Cumulative offer mechanisms (COMs) have received special attention in the literature. In an COM, individuals propose contracts according to a strict ordering $\triangleright$ of the elements of $X$. In each step, some individual who does not currently have a contract held by any institution proposes his most preferred contract that has not yet been proposed. Then, each institution chooses its most preferred set of contracts according to its choice rule and holds this set until the next step. When multiple individuals are able to propose in the same step, the individual who proposes is determined by the ordering $\triangleright$.
@@
Some individual $i^{l}\in\ I$, for whom no institution currently holds a contract, proposes her most preferred contact that has not yet been rejected $x^{l}\in X_{i^{l}}\setminus\left(\bigcup_{j\in S}A_{j}^{l}\right)$. Institution $\mathbf{s}\left(x^{l}\right)$ holds the set of contracts in $C_{\mathbf{s}\left(x^{l}\right)}\left(A_{\mathbf{s}\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}\right)$ and rejects all other contracts in $A_{s\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}$; institutions $s^{\prime}\neq\mathbf{s}\left(x^{l}\right)$ continue to hold all contracts they held at the end of Step $l-$1. Set $A_{\mathbf{s}\left(x^{l}\right)}^{l+1}=A_{\mathbf{s}\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}$, and set $A_{s^{\prime}}^{l+1}=A_{s^{\prime}}^{l}$ for each $s^{\prime}\neq\mathbf{s}\left(x^{l}\right)$.
@@
A direct corollary of Theorem 1 is that the COM is the only candidate if the policy makers insist on stability and strategy-proofness.
@@
The COM is the unique mechanism that is stable and strategy-proof when institutions’ choice rules are in the GS family.
5. Practical Choice Rules in the GS Family
Policy goals in many real-world matching markets naturally give rise to choice rules that are in the GS family. In this section, we present three applications of real-world resource allocation problems where the proposed choice rules that emerge from policy desiderata are in the GS family. Namely, we analyze (i) the aggregate choice rule in which the meritorious horizontal choice rule selects candidates in each category in allocation government jobs in India (Sönmez and Yenmez (2022) and (ii) the simultaneous open-reserve open-choice rules proposed for high school admissions in China (Hu and Zhang (2025)).

### v3

@@ 2. Preliminaries
We consider a centralized many-to-one matching market. There is a finite set of individuals $I=\{i_{1},...,i_{n}\}$ and a finite set of institutions $S=\{s_{1},...,s_{m}\}$. We let $\overline{q}_{s}$ be the capacity of the institution $s$.
@@
There is a finite set of contracts $X$. Each contract $x\in X$ is between an individual $\mathbf{i}(x)$ and an institution $\mathbf{s}(x)$. For any set of contracts $Y\subseteq X$, let $\mathbf{i}(Y)\equiv\bigcup_{x\in Y}\{\mathbf{i}(x)\}$ be the set of individuals who have a contract associated with them in $Y$. Similarly, we let $\mathbf{s}(Y)\equiv\bigcup_{x\in Y}\{\mathbf{s}(x)\}$. We also define $Y_{i}\equiv\{x\in Y\mid i=\mathbf{i}(x)\}$ for each $i\in I$ and $Y_{s}\equiv\{x\in Y\mid s=\mathbf{s}(x)\}$ for each $s\in S$.
@@
Each individual $i\in I$ has a strict preference $P_{i}$ over contracts in $X_{i}=\{x\in X|\mathbf{i}(x)=i\}$ and an outside option $\emptyset$. We denote $\mathcal{P}_{i}$ the set of all strict preferences individual $i$ may have. A contract $x\in X_{i}$ is acceptable to $i$ with respect to $P_{i}$ if $xP_{i}\emptyset$. We let $R_{i}$ be the weak preference relation associated with $P_{i}$. $xR_{i}y$ if and only if $xP_{i}y$ or $x=y$.
@@
Each institution $s\in S$ has multiunit demand and is endowed with an aggregate choice rule $C_{s}$ that describes how $s$ would choose from any set of contracts. For all $Y\subseteq\ X$ and $s\in S$, the aggregate choice rule $C_{s}$ is such that $C_{s}(Y)\subseteq Y_{s}$ and $\mid C_{s}(Y)\mid\leq\overline{q}_{s}$. In the next section, we introduce a rich class of choice rules for institutions.
A matching is a set of contracts $Y\subseteq X$. A matching $Y\subseteq X$ is feasible if $\mid Y_{i}\mid\leq 1$ for all $i\in I$. A feasible matching $Y\subseteq X$ is stable if
@@
Condition (1) is the standard individual rationality condition for individuals. Condition (2) requires that institutions’ selection procedures be respected and can be interpreted as an individual rationality condition for institutions. [^5] This condition ensures that criteria, such as diversity and affirmative action, that are encoded in institutions’ choice rules are complied with; See alva2023choice for a detailed discussion of this property. Condition (3) is the standard no blocking condition and requires that there be no set of contracts $Z$ such that all institutions and individuals associated with contracts in $Z$ prefer to receive the contracts in $Z$.
Stability and strategy-proofness are fundamental objectives in market design. Stability precludes scenarios in which agents can profitably rearrange their matches after the market has cleared (gale1962college). Stability is considered a crucial determinant of the success of a centralized clearinghouse (roth1990new). Stability also ensures that the institutional objectives, typically set by policymakers and encoded in institutions’ choice rules, are satisfied. Moreover, there is a sense in which stability implies fairness (or no justified envy). [^6] See romm2024stability for a detailed analysis of the relationship between stability and no justified envy in a matching model with contracts. Strategy-proofness establishes truthful preference revelation as a dominant strategy, eliminating incentives for strategic misrepresentation, and thus simplifying market participation while ensuring that allocations derive from accurate preference data.
Given a profile of aggregate choice rules of institutions $C=(C_{s})_{s\in\ S}$, a mechanism $\psi(\cdot;C)$ maps preference profiles $P=(P_{i})_{i\in\ I}$ to matchings. Unless otherwise stated, we assume that institutions’ choice rules are fixed. A mechanism $\psi$ is stable if the outcome of $\psi$, $\psi(P;C)$, is a stable matching for every preference profile $P$, given a profile of the institutions’ choice rules $C$. A mechanism $\psi$ is strategy-proof if, given a profile of institutions’ choice rules $C$, for every preference profile $P$ and for each individual $i\in I$, there is no $\widetilde{P}_{i}\in\mathcal{P}_{i}$ such that
@@
3. Generalized Sequential Choice Rules
This section introduces a class of choice functions, named Generalized Sequential (GSq). Consider an institution $s$ that has a set of divisions $\mathcal{K}_{s}=\{1,...,K_{s}\}$. Each division $k\in\mathcal{K}_{s}$ has an associated choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ that specifies the contracts that division $k$ chooses given a set of contracts and a capacity. The capacity of division $k>1$ may vary as a function of the number of vacant positions in the preceding divisions. We require that for a set of contracts $Y\subseteq\ X_{s}$ and a capacity of $\lambda\in\mathbb{Z}_{+}$, $\mid C_{s}^{k}(Y;\lambda)\mid\leq\lambda$ and that
$x,x^{{}^{\prime}}\in C^{k}_{s}(Y;\lambda)\implies\mathbf{i}(x)\neq\mathbf{i}(x^{{}^{\prime}}).$
That is, the choice rule of a division can select at most one contract for an individual.
Given a set of contracts $Y\equiv Y^{1}\subseteq\ X$ , a capacity $\overline{q}_{s}$ for institution $s$, and a capacity $\overline{q}_{s}^{1}$ for division 1, we compute the set chosen $C_{s}(Y,\overline{q}_{s})$ sequentially in $K_{s}$ steps, where division $k$ chooses in step $k$, for $k=1,...,K_{s}$, as follows:
Step 1
$r_{1}=\overline{q}_{s}^{1}-\mid C_{s}^{1}(Y^{1};\overline{q}_{s}^{1})\mid$
be the number of vacant positions in division 1. Moreover, let
Step k ($2\leq k\leq K_{s})$
The capacity of division $k$ is considered as a function of the number of unfilled positions in the preceding divisions, $q_{s}^{k}(r_{1},...,r_{k-1})$, since we allow redistribution of vacant positions in the preceding divisions. [^7] We give the details of capacity redistributions in the next subsection, titled transfer policy.
be the number of unfilled positions in division $k$. Moreover, let
The union of divisions’ choices is "the" institution’s chosen set. That is,
Transfer Policy
We impose the following mild condition on transfer functions.
Definition 1.
$q_{s}^{j}(\widetilde{r}_{1},...,\widetilde{r}_{j-1})\geq q_{s}^{j}(r_{1},...,r_{j-1});$ (1)
and
$\sum_{m=2}^{j}[q_{s}^{m}(\tilde{r}_{1},...,\widetilde{r}_{m-1})-q_{s}^{m}(r_{1},...,r_{m-1})]\leq\sum_{m=1}^{j-1}[\widetilde{r}_{m}-r_{m}].$ (2)
The first condition simply says that the capacity of a division is a (weakly) increasing function of the number of vacancies in the preceding divisions. The second one states that no extra slot is created, in the sense that any increase in the capacity of a division must come from slots left vacant by preceding divisions.888westkamp2013analysis introduced the first and second conditions in a simpler matching model without contracts as different properties. westkamp2013analysis called the first one ”monotonicity” and the second one non-excessive reduction” property. We adopted these properties to a more general matching model with contracts and combined them in a single property, called monotonicity, to reduce the number of conditions dealt with.
Conditions on Divisions’ Choice Rules
A choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ satisfies the substitutes property if, for all $q\in\mathbb{Z}_{+}$ and for all $x,y\in X$ and $Y\subseteq X\setminus\{x,y\}$,
$\mid C_{s}^{k}(Y;q)\mid\leq\mid C_{s}^{k}(Y\cup\{x\};q)\mid.$
The substitutes property and size monotonicity jointly imply the irrelevance of rejected contracts (IRC) condition (aygun2013matching). A choice rule $C_{s}^{k}$ satisfies the IRC condition if, for any $Y\subseteq X$ and any $x\in X\setminus Y$,
Both the substitutes property and size monotonicity are standard conditions that impose restrictions on the choice function when the set of alternatives changes, while the capacity remains unchanged. The next property imposes conditions on the choice function when the capacity increases while the set of alternatives remains unchanged. This property is in the spirit of the expansion property of the choice rules defined in chambers2017choice.
Definition 2.
$C_{s}^{k}(Y,q)\subseteq C_{s}^{k}(Y,q+\Delta),\$(3)
and
$\mid C_{s}^{k}(Y,q+\Delta)\mid-\mid C_{s}^{k}(Y,q)\mid\leq\Delta.$ (4)
Condition (3) says that if there is an increase in capacity, the choice rule selects every contract chosen before the increase in capacity from any given set of contracts. Condition (4) simply states that if the capacity increases by $\Delta$, then the difference between the number of contracts chosen with the increase in capacity and the initial capacity cannot exceed $\Delta$. The first condition is the expansion property of chambers2017choice. The second condition restricts the magnitude of expansion when capacity is changed.
Definition 3.
An aggregate choice rule $C_{s}$ is a GSq rule if it can be represented as a pair $\left(\left(C_{s}^{k}\right)_{k\in\mathcal{K}_{s}},q_{s}\right)$ of a collection of division choice rules and transfer policy, where each division’s choice rule satisfies the substitutes property, size monotonicity and quota monotonicity, and the transfer policy satisfies monotonicity.
The family of GSq choice rules is a parametrized family of aggregate choice rules derived from choice functions that satisfy both the substitutes property and size monotonicity. The parameterization comes from two primary sources. The transfer function provides one dimension of parameterization. Different monotonic transfer functions generate distinct aggregate choice rules within this family.
Moreover, parameterization arises from the expansion of divisions’ choice rules while maintaining quota monotonicity. A division may be characterized by a sequence of choice rules, with each rule corresponding to a specific integer capacity. By defining the choice rule of division $k$ as $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$, we emphasize that capacity of division $k$ may vary. In a sense, this definition assigns a series of choice functions to division $k$, one for each given integer capacity. When the capacity of a substitutable and size monotonic choice rule is adjusted, the result is a new choice rule that also satisfies these properties. The quota monotonicity property establishes the precise relationship between these capacity-differentiated choice rules.
A substitutable and size monotonic choice rule can be "expanded" to another such rule through multiple different ways in the sense that multiple different choice rules may expand the same choice rule that satisfies the subsitutes property and size monotonicity. This fact further enriches the parametric space of the GSq family.
When each division’s choice rule is q-responsive, which implies the substitutes property, size monotonicity, and QM, the aggregate choice rule becomes a dynamic reserves choice rule introduced in aygun2020dynamic. Therefore, the family of GSq rules nests the family of dynamic reserves choice rules. An important example of a choice rule that satisfies the substitute property, size monotonicity, and QM but violates responsiveness is the meritorious horizontal choice rule introduced in sonmez2022affirmative. The GSq family also nests the family of slot-specific priorities choice rules, introduced by kominers2016matching, because aygun2020dynamic shows that each slot-specific priorities choice rule can be represented as a dynamic reserves choice rule.
The aggregate choice rule in the GSq family may fail to satisfy the substitute property and size monotonicity, even though the divisions’ choice rules satisfy them. However, a stable and strategy-proof mechanism design is still possible.
4. Stable and strategy-proof mechanisms
@@
hatfield2019hidden introduce conditions on the choice rules of the institutions to guarantee a stable and strategy-proof matching mechanism.
Definition 4 (hatfield2019hidden).
@@
Let $C_{s}(\cdot;q_{s})$ be an GSq choice rule induced by the transfer function $q_{s}$. We define a related choice function $\overline{C}_{s}(\cdot;q_{s})$ as follows:
@@
Given $Y^{1}$ and $\overline{q}^{1}_{s}$, division 1 chooses $C^{1}_{s}(Y^{1};\overline{q}^{1}_{s})$. Let $r_{1}=\overline{q}^{1}_{s}-\mid C^{1}_{s}(Y^{1};\overline{q}^{1}_{s})\mid$. Let $Y^{2}\equiv Y^{1}\setminus C^{1}_{s}\left(Y^{1};\overline{q}^{1}_{s}\right)$.
@@
Given $Y^{k}$ and its capacity $q^{k}_{s}(r_{1},...,r_{k-1})$, division $k$ chooses $C^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))$. Let $r_{k}=q^{k}_{s}(r_{1},...,r_{k-1})-\mid C^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))\mid$. Let $Y^{k+1}\equiv Y^{k}\setminus C^{k}_{s}\left(Y^{k};q^{k}_{s}\left(r_{1},\ldots,r_{k-1}\right)\right)$.
The union of the’ chosen sets of the divisions is "the" institution’s chosen set. That is,
$\overline{C}_{s}(Y;q_{s})\equiv C^{1}_{s}(Y^{1};\overline{q}^{1}_{s})\cup\bigcup_{k=2}^{K_{s}}{C^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))}.$
The difference between $C_{s}$ and $\overline{C}_{s}$ is as follows: In the computation of $C_{s}$, if a contract of an individual is chosen by some division, then her other contracts are removed for the rest of the procedure. According to $\overline{C}_{s}$, if an individual’s contract is chosen by division $k$, then her other contracts will still be available for the remaining divisions. Note that the completion uses the same divisional choice rules, i.e. $\bar{C}_{s}^{k}=C_{s}^{k}$, for all $k=1,...,K_{s}$.
We are now ready to state our main result.
@@
Cumulative offer mechanisms (COMs) have received special attention in the literature. In an COM, individuals propose contracts according to a strict ordering $\triangleright$ of the elements of $X$. In each step, some individual who does not currently have a contract held by any institution proposes his most preferred contract that has not yet been proposed. Then, each institution chooses its most preferred set of contracts according to its choice rule and holds this set until the next step. When multiple individuals are able to propose in the same step, the individual who proposes is determined by ordering $\triangleright$.
@@
Some individual $i^{l}\in\ I$, for whom no institution currently holds a contract, proposes her most preferred contact that has not yet been rejected $x^{l}\in X_{i^{l}}\setminus\left(\bigcup_{j\in S}A_{j}^{l}\right)$. Institution $\mathbf{s}\left(x^{l}\right)$ holds the set of contracts in $C_{\mathbf{s}\left(x^{l}\right)}\left(A_{\mathbf{s}\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}\right)$ and rejects all other contracts in $A_{s\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}$; institutions $s^{\prime}\neq\mathbf{s}\left(x^{l}\right)$ continue to hold all contracts they held at the end of Step ll-1. Set $A_{\mathbf{s}\left(x^{l}\right)}^{l+1}=A_{\mathbf{s}\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}$, and set $A_{s^{\prime}}^{l+1}=A_{s^{\prime}}^{l}$ for each $s^{\prime}\neq\mathbf{s}\left(x^{l}\right)$.
@@
The order of contracts offered during the COM does not matter. The outcome of the COM is independent of the order of proposals (hatfield2017restud). [^9] In their Proposition 3, hatfield2017restud show that observable substitutability of institutions’ choice rules is sufficient for the order independence of the COM. Choice rules in the GSq family satisfy observable substitutability.
@@
hatfield2017restud show that when institutions’ choice rules satisfy three conditions –observable substituablity, observable size monotonicity, and not manipulability via contractual terms – the COM is the unique mechanism that is stable and strategy-proof. Moreover, hatfield2017restud show that choice rules that admit completions that satisfy the substitute property and size monotonicity satisfy these three conditions. Therefore, combining these results with our Theorem 1, we obtain the following important corollary.
@@
The COM is the unique mechanism that is stable and strategy-proof when institutions’ choice rules are in the GSq family.
5. Practical Choice Rules in the GSq Family
Policy goals in many real-world matching markets naturally give rise to choice rules that are in the GSq family. In this section, we present three applications of real-world resource allocation problems where the proposed choice rules that emerge from policy desiderata are in the GSq family. Namely, we analyze (i) the aggregate choice rule in which the meritorious horizontal choice rule selects candidates in each category in allocation government jobs in India (sonmez2022affirmative) and (ii) the simultaneous open-reserve open-choice rules proposed for high school admissions in China (hu2025verifiable).